I need to find all three sides of a right triangle (use the pythagorean therum to make a quadratic equation then solve)
Side A-X
Side B-X+1
Side C-X+2

Answer by ewatrrr(24785)

 
Hi,
Right Triangle with Sides x, (x+1) and (x+2)...
Applying the Pythagorean Theorem
x^2 + (x+1)^2 = (x+2)^2
x^2 + x^2 + 2x + 1 = x^2 + 4x + 4
x^2 -2x - 3 = 0
factoring
(x-3)(x+1) = 0
x = 3 (throwing out negative solution -1 for length)
Sides are 3, 4, 5